# Vulnerabilities of smart contracts and mitigation schemes: A Comprehensive Survey

### v1

@@ III-C3 Protection Measures
In order to avoid integer overflows and underflows, arithmetic operations should be properly implemented by checking the operators and operands before performing the operation. It is recommended to use the assert() and require() modifiers. It is also advisable to use the library for arithmetic functions, called SafeMath by OpenZeppelin [18]. Solidity version 0.8.0 has included this library; if an overflow/underflow occurs, the transaction will revert.A prevalent method for detecting this type of vulnerability involves the use of taint analysis, a technique that we will elaborate on in Section IV.
III-C4 Example
@@
Listing 7 showcases a smart contract managing fractional tokens of an NFT, vulnerable to underflow. If an attempt is made to transfer more tokens than available, under Solidity versions prior to 0.8.0, this could lead to an underflow, setting the token balance to an incorrect high value, potentially enabling malicious tokens distribution.
@@
Given Solidity 0.8.0’s automatic checks for arithmetic operations, the primary vulnerability of underflows and overflows is directly addressed by the compiler. However, we can implement additional logic to further safeguard the integrity of fractional tokens calculations and transfers. An improved version of the smart contract is presented in Listing 8.
III-D Mishandled Exceptions
III-D1 Description
@@
Solidity provides two primary paradigms for interacting with external contracts: direct contract calls and low-level calls. Direct contract calls involve invoking functions directly on known contract interfaces, which inherently revert on failure, thereby throwing an exception. Conversely, low-level calls, executed via methods like call(), delegatecall(), and callcode(), return a boolean success flag instead of reverting on exceptions [21]. These calls do not inherently revert transaction execution upon failure; instead, they return false, necessitating explicit checks of their return values to ensure intended execution flow. Failure to adequately check the result of a low-level call can lead to unintended execution continuation, potentially compromising contract logic and security. This vulnerability was notably exploited in the King of Ether game, where the smart contract’s failure to verify the result of a send() operation led to discrepancies in payments, resulting in users overpaying or underpaying [22].
III-D2 Implications on NFT Fractionalization
@@ III-D3 Example
III-E Code Injection via delegatecall
III-E1 Description
@@
There is a unique type of method called delegate call. The DELEGATECALL opcode is similar to a conventional message call with the exception that the code performed at the target address is executed in the context of the calling contract. The current address, storage, and balance refer to the calling contract. This enables a smart contract to load code from another smart contract dynamically at runtime. Calling into untrusted contracts via delegatecall() is particularly risky since the code at the target smart contract has complete control over the caller’s balance; thus, it can modify any of the caller’s storage data.
III-E2 Implications on NFT Fractionalization
@@
Malicious code executed through delegatecall might tamper with the contract’s ownership or control mechanisms, enabling attackers to reroute assets or funds. Furthermore, attackers could change the revenue distribution logic, redirecting profits meant for rightful owners to unauthorized entities. In extreme cases, like the second Parity multi-sig attack where an attacker took over three main Parity wallets and stole $31 million [23] , these vulnerabilities could let attackers completely take over the NFT fractionalization contract. This means they could push out the real owners and managers, potentially locking, freezing, or stealing assets.
III-E3 Protection Measures
@@
When employing delegatecall(), caution is paramount, especially when dealing with contracts not fully trusted. It’s critical to avoid making delegatecall() to addresses derived from user inputs unless there’s a rigorous verification process against a list of trusted contracts. This precaution helps in ensuring that only known, secure contracts can be interacted with, significantly reducing the risk of malicious interference.
@@
Solidity’s library keyword facilitates the creation of library contracts, designed to be stateless and immune to being destroyed. By defining a contract as a stateless library, it’s guaranteed that the executing code cannot alter the storage data of the calling contract. This design principle is crucial for minimizing risks associated with storage context issues that delegatecall() might introduce. Ensuring that delegatecall() is only used with contracts declared as libraries is a solid practice.
III-E4 Example
@@
Listing 11 demonstrates how an NFT fractionalization contract might be exposed to a code injection vulnerability through unsafe use of delegatecall. In this contract, executeDistributionLogic can execute arbitrary logic via delegatecall, based on the bytecode provided in data. If logicContract points to a malicious contract, it could lead to unintended alterations in the contract’s state, including token ownership manipulation.
@@
To mitigate the risk of code injection via delegatecall, the contract should strictly control the update of the logicContract address and ensure that only verified, safe operations are executable. Also, we introduce a whitelist of approved logic contract addresses and require that any updates to logicContract come from this whitelist. Listing 12 illustrates this changes.
III-F Randomness Using Block Information
III-F1 Description
@@
In blockchain-based applications, certain functionalities might require randomness—for instance, distributing a rare NFT fractionally among participants or deciding the winner of a unique NFT in a lottery, block information, such as block hash and block timestamp, can be used to achieve this. This information, however, may be anticipated and slightly modified by miners. When the block timestamp is used as the trigger condition to execute the transaction, it creates a vulnerable situation; indeed, dishonest miners can exploit the value of the block timestamp in an unethical manner. This vulnerability was exploited in GovernMental, a Ponzi scheme game [24]. The player who joined the round last and stayed for at least a minute was compensated according to the game rules. A miner who is also a player, might change the timestamp to make it look like he/she was the last to join for more than a minute and therefore collect the reward.
III-F2 Implications on NFT Fractionalization
@@
Randomness Using Block Information could impact Fractionalization solutions. For instance, if randomness derived from block attributes decides the allocation of rare NFT fractions or the winners of rewards, it could lead to outcomes unfairly attribute in favor of those with the ability to influence block information. Similarly, auctions for selling fractionalized NFTs could be manipulated, allowing miners or others with insider advantages to affect the auction’s outcome by adjusting the voting period for example.
III-F3 Protection Measures
@@
To address the vulnerabilities associated with using block information for generating randomness in smart contracts, adopting more reliable sources of randomness is essential. One solution is employing oracles [25] or other external sources that provide verified randomness is a robust solution.
@@
Alternatively, cryptographic commitment schemes [26], exemplified by solutions like RANDOA [27] , represent another effective approach. These schemes involve participants committing to their inputs in a concealed manner, which are later revealed collectively to generate a random outcome. This method ensures that no individual can influence the result based on other participants’ commitments, fostering fairness and security.
III-F4 Example
@@
To mitigate this vulnerability, the platform can utilize an external oracle to provide a source of verified randomness. This approach ensures that the randomness used for fraction distribution is not predictable or manipulable by miners. Contract illustrated in Listing 14 inherits from VRFConsumerBase, provided by Chainlink [28], to securely request and receive verified random numbers. It uses the Chainlink VRF (Verifiable Random Function) to ensure that the randomness determining the distribution of NFT fractions cannot be influenced by participants, including miners.
III-G Other Vulnerabilities & Summary
Table II present the papers that include each vulnerability. We notice that the most mentioned vulnerability is reeantancy, whereas code injection is the least discussed.
Reentrancy | Front Running | Arithmetic | Mishandled exceptions | Code Injection | Randomness | |
[29] | X | |||||
[30] | X | X | X | |||
[31] | X | X | X | X | ||
[32] | X | X | X | |||
[33] | X | |||||
[34] | X | X | X | X | ||
[24] | X | X | X | X | X | |
[35] | X | X | X | X | ||
[36] | X | X | X | X | ||
[37] | X | X | X | X | ||
[5] | X | X | X | X | ||
[38] | X | |||||
[6] | X | |||||
[18] | X | X | X | |||
[39] | X | X | X | |||
[40] | X | X | ||||
[41] | X | |||||
[42] | X | |||||
[43] | X |
The Smart Contract Weakness Classification Registry (SWC) [44] is a community database of known smart contract vulnerabilities; it is often up to date. For each issue, it includes a description, code samples, and protection measures. Currently, the registry contains 36 vulnerabilities. One has to check this registry, often, to stay up to speed on the newest threats.
@@ IV Detection methods
In this section, we present the most common approaches for detecting smart contract vulnerabilities. They include static analysis, dynamic analysis, and formal verification. We will briefly describe each of these approaches and identify their benefits and limitations.
IV-A Static Analysis
@@
Static code analysis is a technique of debugging that involves reviewing source code prior to running it; this technique is also known as white-box testing [46]. It is accomplished by comparing a set of code to a set of predefined coding rules. In fact, there are a number of techniques to examine static source code; they can be incorporated into a single solution. Compiler technologies are frequently used to develop these techniques such as Taint Analysis and Data Flow Analysis [47]. In this Section we present some static analysis techniques, namely Control Flow Graph, Taint Analysis, and Symbolic Analysis.
IV-A1 Control Flow Graph
@@
The Control Flow Graph (CFG) is a directed graph that describes the flow of control. The flow of execution of source code A graph node denotes a basic block with no jumps; directed edges represent jumps from one block to another. If a node only has an exit edge, it is referred to as an entry. And if it has only an entry edge, it is referred to as an exit block. In Figure 15, Block 1448 is an entry block, and blocks 1451 and 1452 are exit blocks.
@@
Figure 16 shows the common process to create CFG in the context of Ethereum. The first step is Parsing bytecode. Typically, the algorithm will extract the compiler version from the metadata; then, it will parse the remaining bytecode (without the metadata) into opcodes [48]. The parsing stage is straightforward because each of the bytecode’s two characters represent one opcode (for example 0x04 represents DIV). The complete list of opcodes can be found in the Ethereum yellow paper [49]. The second step is Identification of basic blocks. A basic block is a set of opcodes that run in succession between a jump target and a jump instruction, with no other instructions interrupting the control flow. To determine the basic blocks, we need to identify the opcodes that alter the execution of the control flow; some examples are presented in Table III. The JUMPDEST instruction starts a new basic block, whereas the other opcodes terminate blocks. Each block is identified by an offset that represents the position of its first opcode in the bytecode. The last step will be computing the edge by checking the destination offset; this step is not trivial as the destination is not an opcode parameter. In the literature, there are two types of jumps: (a) A pushed jump is a JUMP followed by a PUSH opcode, allowing the target offset to be determined simply by glancing at the data in the previous PUSH opcode; and (b) Orphan jump is not followed by a PUSH; indeed, its destination is not computed directly [50]. The computation of orphan jumps depends on each implementation and it is critical for the completeness of the graph.
Operation | Opcode | Description |
Flow | JUMP | Alters the program counter |
Alters the program counter | ||
Operations | JUMPI | Conditionally alters the program counter |
JUMPDEST | Marks a valid destination for jumps | |
System | STOP | Halts execution |
operations | REVERT | Halts execution reverting state changes |
but returning data and remaining gas | ||
RETURN | Halts execution returning output data | |
INVALID | Designated invalid instruction | |
SELFDESTRUCT | Halts execution and registers account | |
@@ IV-A2 Taint Analysis
Taint Analysis aims to detect input variables where data comes from an untrusted source that could be controlled by an attacker (e.g. environmental data or function parameters). These data inputs are tainted by the taint analysis tool and then traced back to potentially sensitive functions, such as security checks or storage access, often know as sinks. If the tainted variable flows into a sink, it is marked as a vulnerability [41].
@@
In the context of Ethereum smart contracts, this technique is usually used after generating CFG. It exploits opcode that an attacker may employ to insert data into the control flow; it can be divided into block information, environmental information, and flow information. Table IV shows an example of opcode that allows data insertion can be found in .
Operation | Opcode | Description |
Block | GASLIMIT | Gets the current block gas limit |
information | TIMESTAMP | Gets the current block timestamp |
Environment | CALLER | Gets caller address |
information | CALLDATALOAD | Gets input data of current |
environment | ||
CALLDATACOPY | Copies input data in current | |
environment to memory | ||
Flow | SLOAD | Loads word from storage |
Operations | MLOAD | Loads word from memory |
@@
An example of taint analysis is to detect arithmetic bugs. The first step is to define the possible entrusted inputs (source) such as CALLDATALOAD and SLOAD. The affected memory, storage or stack location is tagged. The next step is to check whether the flow of tagged input contains arithmetic operations such as ADD, MUL and SUB with no catch mechanism (sink). For example, a catch mechanism is reverted in case of DIV with 0. If the tagged input falls into a sink, it will be marked as a possible vulnerability.
IV-A3 Symbolic Analysis
@@
Symbolic execution is a way of abstracting the execution of a program such that it can span numerous pathways through the code. The program is run with symbols as inputs, and the outputs are expressed in the form of the symbolic inputs. Each symbolic path has a condition; it is a formula created by collecting constraints that must be satisfied by those inputs in order for the execution to continue on that path. If the condition is unsatisfiable, the path is infeasible; otherwise, the path is feasible [51]. Symbolic analysis is usually paired with CFGs. In order to find exploits, most tools employ the Satisfiability Modulo Theories (SMT) [52] solver to check whether the symbolic output is feasible or provide a counterexample if it is not.
IV-B Dynamic Analysis

### v2

@@ 3.3.3 Protection Measures
Arithmetic operations should be appropriately implemented by checking the operators and operands before operating to avoid integer overflows and underflows. It is recommended to use the assert() and require() modifiers. Using the library for arithmetic functions is also advisable, called SafeMath by OpenZeppelin [18]. Solidity version 0.8.0 has included this library, so the transaction will revert if an overflow/underflow occurs. A prevalent method for detecting this type of vulnerability involves taint analysis, a technique we will elaborate on in Section 4.
3.3.4 Example
@@
Listing 7 showcases a smart contract managing fractional tokens of an NFT, vulnerable to underflow. If an attempt is made to transfer more tokens than available under Solidity versions before 0.8.0, this could lead to an underflow, setting the token balance to an incorrect high value and potentially enabling malicious token distribution.
@@
Given Solidity 0.8.0’s automatic checks for arithmetic operations, the compiler directly addresses the primary vulnerability of underflows and overflows. However, we can implement additional logic to further safeguard the integrity of fractional token calculations and transfers. An improved version of the smart contract is presented in Listing 8.
3.4 Mishandled Exceptions
3.4.1 Description
@@
Solidity provides two primary paradigms for interacting with external contracts: direct contract calls and low-level calls. Direct contract calls involve invoking functions directly on known contract interfaces, which inherently revert to failure, thereby throwing an exception. Conversely, low-level calls, executed via methods like call(), delegatecall(), and callcode(), return a boolean success flag instead of reverting on exceptions [21]. These calls do not inherently revert transaction execution upon failure; instead, they return false, necessitating explicit checks of their return values to ensure the intended execution flow. Failure to adequately check the result of a low-level call can lead to unintended execution continuation, potentially compromising contract logic and security. This vulnerability was notably exploited in the King of Ether game, where the smart contract’s failure to verify the result of a send() operation led to discrepancies in payments, resulting in users overpaying or underpaying [22].
3.4.2 Implications on NFT Fractionalization
@@ 3.4.3 Example
3.5 Code Injection via delegatecall
3.5.1 Description
@@
There is a unique method called a delegate call. The DELEGATECALL opcode is similar to a conventional message call, except that the code performed at the target address is executed in the context of the calling contract. The current address, storage, and balance refer to the calling contract. This dynamically enables a smart contract to load code from another smart contract at runtime. Calling into untrusted contracts via delegatecall() is particularly risky since the code at the target smart contract has complete control over the caller’s balance; thus, it can modify any of the caller’s storage data.
3.5.2 Implications on NFT Fractionalization
@@
Malicious code executed through delegatecall might tamper with the contract’s ownership or control mechanisms, enabling attackers to reroute assets or funds. Furthermore, attackers could change the revenue distribution logic, redirecting profits meant for rightful owners to unauthorized entities. In extreme cases, like the second Parity multi-sig attack where an attacker took over three main Parity wallets and stole $31 million [23], these vulnerabilities could let attackers take over the NFT fractionalization contract entirely. They could push out the real owners and managers, potentially locking, freezing, or stealing assets.
3.5.3 Protection Measures
@@
When employing delegatecall(), caution is paramount, especially when dealing with contracts not fully trusted. It’s critical to avoid making delegatecall() to addresses derived from user inputs unless there’s a rigorous verification process against a list of trusted contracts. This precaution helps ensure that only known, secure contracts can be interacted with, significantly reducing the risk of malicious interference.
@@
Solidity’s library keyword facilitates the creation of library contracts designed to be stateless and immune to destruction. By defining a contract as a stateless library, it’s guaranteed that the executing code cannot alter the storage data of the calling contract. This design principle is crucial for minimizing risks associated with storage context issues that delegatecall() might introduce. Ensuring that delegatecall() is only used with contracts declared as libraries is a solid practice.
3.5.4 Example
@@
Listing 11 demonstrates how an NFT fractionalization contract might be exposed to a code injection vulnerability through unsafe use of delegatecall. In this contract, executeDistributionLogic can execute arbitrary logic via delegatecall based on the bytecode provided in the data. If logicContract points to a malicious contract, it could lead to unintended alterations in its state, including token ownership manipulation.
@@
To mitigate the risk of code injection via delegatecall, the contract should strictly control the update of the logicContract address and ensure that only verified, safe operations are executable. Also, we introduce a whitelist of approved logic contract addresses and require that any updates to logicContract come from this whitelist. Listing 12 illustrates these changes.
3.6 Randomness Using Block Information
3.6.1 Description
@@
In blockchain-based applications, certain functionalities might require randomness—for instance, distributing a rare NFT fractionally among participants or deciding the winner of a unique NFT in a lottery; block information, such as block hash and block timestamp, can be used to achieve this. This information, however, may be anticipated and slightly modified by miners. When the block timestamp is used as the trigger condition to execute the transaction, it creates a vulnerable situation; dishonest miners can exploit the value of the block timestamp in an unethical manner. This vulnerability was exploited in GovernMental, a Ponzi scheme game [24]. The player who joined the round last and stayed for at least a minute was compensated according to the game rules. A miner who is also a player might change the timestamp to make it look like they were the last to join for more than a minute and, therefore, collect the reward.
3.6.2 Implications on NFT Fractionalization
@@
Randomness Using Block Information could impact Fractionalization solutions. For instance, if randomness derived from block attributes decides the allocation of rare NFT fractions or the winners of rewards, it could lead to outcomes unfairly attributed in favor of those with the ability to influence block information. Similarly, auctions for selling fractionalized NFTs could be manipulated, allowing miners or others with insider advantages to affect the auction’s outcome by adjusting the voting period, for example.
3.6.3 Protection Measures
@@
Adopting more reliable sources of randomness is essential to address the vulnerabilities associated with using block information to generate randomness in smart contracts. One solution is employing oracles [25] or other external sources that provide verified randomness robust solution.
@@
Alternatively, cryptographic commitment schemes [26], exemplified by solutions like RANDOA [27], represent another practical approach. These schemes involve participants committing to their inputs in a concealed manner, which are later revealed collectively to generate a random outcome. This method ensures no individual can influence the result based on other participants’ commitments, fostering fairness and security.
3.6.4 Example
@@
To mitigate this vulnerability, the platform can utilize an external oracle to provide a source of verified randomness. This approach ensures that the randomness used for fraction distribution is not predictable or manipulable by miners. Contract illustrated in Listing 14 inherits from VRFConsumerBase, provided by Chainlink [28], to securely request and receive verified random numbers. It uses the Chainlink VRF (Verifiable Random Function) to ensure that participants, including miners, cannot influence randomness in determining the distribution of NFT fractions.
3.7 Other Vulnerabilities & Summary
Table II presents the papers that include each vulnerability. We notice that the most mentioned vulnerability is reentrancy, whereas code injection is the least discussed.
Reentrancy |
Front Running |
Arithmetic |
Mishandled exceptions |
Code Injection |
Randomness |
| [29] |
X |
| [30] |
X |
X |
X |
| [31] |
X |
X |
X |
X |
| [32] |
X |
X |
X |
| [33] |
X |
| [34] |
X |
X |
X |
X |
| [24] |
X |
X |
X |
X |
X |
| [35] |
X |
X |
X |
X |
| [36] |
X |
X |
X |
X |
| [37] |
X |
X |
X |
X |
| [5] |
X |
X |
X |
X |
| [38] |
X |
| [6] |
X |
| [18] |
X |
X |
X |
| [39] |
X |
X |
X |
| [40] |
X |
X |
| [41] |
X |
| [42] |
X |
| [43] |
X |
The Smart Contract Weakness Classification Registry (SWC) [44] is a community database of known smart contract vulnerabilities; it is often up to date. Each issue includes a description, code samples, and protection measures. Currently, the registry contains 36 vulnerabilities. One has to check this registry often to stay up to speed on the newest threats.
@@ 4 Detection methods
In this section, we present the most common approaches for detecting smart contract vulnerabilities. These include static analysis, dynamic analysis, and formal verification. We will briefly describe each approach and identify its benefits and limitations.
4.1 Static Analysis
@@
Static code analysis is a technique of debugging that involves reviewing source code before running it; this technique is also known as white-box testing [46]. It is accomplished by comparing a set of codes to predefined coding rules. There are several techniques to examine static source code; they can be incorporated into a single solution. Compiler technologies are frequently used to develop these techniques, such as Taint Analysis, and Data Flow Analysis [47]. This section presents some static analysis techniques, namely Control Flow Graph, Taint Analysis, and Symbolic Analysis.
4.1.1 Control Flow Graph
@@
The Control Flow Graph (CFG) is a directed graph describing the control flow. The flow of execution of source code A graph node denotes a basic block with no jumps; directed edges represent jumps from one block to another. If a node only has an exit edge, it is referred to as an entry. And if it has only an entry edge, it is referred to as an exit block. In Figure 15, Block 1448 is an entry block, and blocks 1451 and 1452 are exit blocks.
@@
Figure 16 shows the standard process to create CFG in the context of Ethereum. The first step is Parsing bytecode. Typically, the algorithm will extract the compiler version from the metadata; then, it will parse the remaining bytecode (without the metadata) into opcodes [48]. The parsing stage is straightforward because each bytecode’s two characters represent one opcode (0x04 represents DIV). The complete list of opcodes can be found in the Ethereum yellow paper [49]. The second step is Identification of basic blocks. A basic block is a set of opcodes that run in succession between a jump target and a jump instruction, with no other instructions interrupting the control flow. To determine the basic blocks, we need to identify the opcodes that alter the execution of the control flow; some examples are presented in Table III. The JUMPDEST instruction starts a new basic block, whereas the other opcodes terminate blocks. Each block is identified by an offset representing its first opcode’s position in the bytecode. The last step will be computing the edge by checking the destination offset; this step is not trivial as the destination is not an opcode parameter. In the literature, there are two types of jumps: (a) A pushed jump is a JUMP followed by a PUSH opcode, allowing the target offset to be determined simply by glancing at the data in the previous PUSH opcode; and (b) Orphan jump is not followed by a PUSH; indeed, its destination is not computed directly [50]. The computation of orphan jumps depends on each implementation and is critical for the completeness of the graph.
Operation |
Opcode |
Description |
Flow |
JUMP |
Alters the program counter |
Alters the program counter |
Operations |
JUMPI |
Conditionally alters the program counter |
JUMPDEST |
Marks a valid destination for jumps |
System |
STOP |
Halts execution |
operations |
REVERT |
Halts execution reverting state changes |
but returning data and remaining gas |
RETURN |
Halts execution returning output data |
INVALID |
Designated invalid instruction |
SELFDESTRUCT |
Halts execution and registers account |
@@ 4.1.2 Taint Analysis
Taint Analysis aims to detect input variables where data comes from an untrusted source that could be controlled by an attacker (e.g., environmental data or function parameters). These data inputs are tainted by the taint analysis tool and then traced back to potentially sensitive functions, such as security checks or storage access, often known as sinks. If the tainted variable flows into a sink, it is marked as a vulnerability [41].
@@
In the context of Ethereum smart contracts, this technique is usually used after generating CFG. It exploits the opcode that an attacker may employ to insert data into the control flow; it can be divided into block, environmental, and flow information. Table IV shows an example of an opcode that allows data insertion can be found in.
Operation |
Opcode |
Description |
Block |
GASLIMIT |
Gets the current block gas limit |
information |
TIMESTAMP |
Gets the current block timestamp |
Environment |
CALLER |
Gets caller address |
information |
CALLDATALOAD |
Gets input data of current |
environment |
CALLDATACOPY |
Copies input data in current |
environment to memory |
Flow |
SLOAD |
Loads word from storage |
Operations |
MLOAD |
Loads word from memory |
@@
An example of taint analysis is to detect arithmetic bugs. The first step is to define the possible entrusted inputs (source), such as CALLDATALOAD and SLOAD. The affected memory, storage, or stack location is tagged. The next step is to check whether the flow of tagged input contains arithmetic operations such as ADD, MUL, and SUB with no catch mechanism (sink). For example, a catch mechanism is reverted in the case of DIV with 0. If the tagged input falls into a sink, it will be marked as a possible vulnerability.
4.1.3 Symbolic Analysis
@@
Symbolic execution is a way of abstracting the execution of a program such that it can span numerous pathways through the code. The program is run with symbols as inputs, and the outputs are expressed in the form of the symbolic inputs. Each symbolic path has a condition; it is a formula created by collecting constraints that must be satisfied by those inputs for the execution to continue on that path. If the condition is unsatisfiable, the path is infeasible; otherwise, the path is feasible [51]. Symbolic analysis is usually paired with CFGs. To find exploits, most tools employ the Satisfiability Modulo Theories (SMT) [52] solver to check whether the symbolic output is feasible or provide a counterexample if it is not.
4.2 Dynamic Analysis